OpenStudy (anonymous):

Substitution. x = 3 - 2y 2x + 4y = 6

OpenStudy (anonymous):

it is given that x=3-2y for substitution, it is easier for us to make one variable on left hand side and the others on right hand side. x is assigned as 3-2y in the question. we can replace x in 2x+4y=6 with 3-2y like this: 2(3-2y)+4y=6 Do you understand until this step? If yes, I will continue my work to show you the next step

OpenStudy (anonymous):

Yes i am on and understand that step.

hero (hero):

x = 3 - 2y 2x + 4y = 6 simplify the second equation to get x + 2y = 3 then subtract 2y from both sides x = 3 - 2y and realize that the second equation is just the same as the first. and it would be pointless to substitute the same equation into itself.

hero (hero):

But doing so would yield 3 - 2y = 3 - 2y

OpenStudy (anonymous):

2(3-2y)+4y=6 for the bracket, we can multiply every element by the number outside the bracket, like this: 2(3)-2(2y) =6-4y put this back to the original equation 6-4y+4y=6 0=0

OpenStudy (anonymous):

This question cannot be solved because of not sufficient condition given

OpenStudy (anonymous):

The equations given are actually the same.

OpenStudy (anonymous):

For solving two unknowns, we know that we have to get two equations which are not dependent of each other

hero (hero):

I always make sure both equations are simplified first before doing any substitutions.

hero (hero):

Usually, that helps avoid any unnecessary steps or dead ends

hero (hero):

You have to watch out for them, especially with substitutions because sometimes there is no solution or the lines are parallel.